# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

### v1

@@ C.1.1 Case 2
Figure 29: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=4,..,12$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 30: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=13,..,21$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 31: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=22,..,29$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
@@
Figure 32: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{30},\alpha,\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
C.1.2 Fatigue-Crack Size Data
Figure 33: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1...,4$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 34: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=5,..,13$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 35: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=14,..,21$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
@@ C.1.2 Fatigue-Crack Size Data
Figure 36: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\alpha,\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Appendix D Comparison of prediction of error for Parametric and Semi-parametric method
D.1 Case 1
Figure 37: Error produced for predicting residual lifetime. The left and right plots denote prediction error for Bayesian semi-parametric and parametric method for each unit.
D.2 Case 2
Figure 38: Error produced for predicting residual lifetime. The left and right plots denote prediction error for Bayesian semi-parametric and parametric method for each unit.
D.3 Fatigue-Crack Size Data
Figure 39: Error produced for predicting residual lifetime. The left and right plots denote prediction error for Bayesian semi-parametric and parametric method for each unit.

### v2

@@ 8 Conclusion
Figure 14: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by parametric method.
APPENDIX B : MCMC convergence for model parameters
Case 1
@@
Figure 15: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,2,3$, $\alpha$ and $\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Case 2
@@
Figure 16: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,2,3$, $\alpha$ and $\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Case 3
Figure 17: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,2,3$, $\alpha$ and $\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Case 4
Figure 18: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,2,3$, $\alpha$ and $\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Case 5
Figure 19: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,2,3$, $\alpha$ and $\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Fatigue-Crack Size dataset
Figure 20: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,2,3$, $\alpha$ and $\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.